# Are Your Models Still Fair? Fairness Attacks on Graph Neural Networks via Node Injections

### v1

@@ Appendix D Implementation algorithm
Due to space limitation, here we provide the complete training process of NIFA in Algorithm 1 for better understanding. The training process and evaluation process are also literally described in Section 4.4.
@@ Appendix E Additional descriptions on victim models
  • GCN [Hamilton et al., 2017]: Borrowing the concept of convolution from the computer vision domain, GCN employs convolution operation on the graph from a spectral perspective to learn the node embeddings.
  • GraphSAGE [Kipf and Welling, 2017]: Given the potential neighborhood explosion issues in GCN, GraphSAGE samples a fixed number of neighbors at each layer during neighborhood aggregation, which greatly improves the training efficiency.
  • APPNP [Klicpera et al., 2019]: Inspired by PageRank, APPNP decouples the prediction and propagation in the training process, which resolves inherent limited-range issues in message-passing models without introducing any additional parameters.
  • SGC [Wu et al., 2019]: SGC empirically finds the redundancy of non-linear activation function, and achieves comparable performance with much higher efficiency.
  • FairGNN [Dai and Wang, 2021]: Through proposing a sensitive attribute estimator and an adversarial learning module, FairGNN maintains high classification accuracy while reducing unfairness in scenarios with limited sensitive attribute information.
  • FairVGNN [Wang et al., 2022b]: FairVGNN discovers the leakage of sensitive information during information propagation of GNN models, and generates fair node features by automatically identifying and masking sensitive-correlated features.
  • FairSIN [Yang et al., 2024]: Instead of filtering out sensitive-related information for fairness, FairSIN deploys a novel sensitive information neutralization mechanism. Specifically, FairSIN will learn to introduce additional fairness facilitating features (F3) during message propagation to neutralize sensitive information while providing more non-sensitive information.
@@ Appendix F Additional descriptions on baselines
  • AFGSM [Wang et al., 2020]: As a node injection-based attack, AFGSM designs an approximation strategy to linearize the victim model and then generates adversarial perturbation efficiently. In general, AFGSM is scalable to much larger graphs.
  • TDGIA [Zou et al., 2021]: Aiming at attacking the model performance, TDGIA first introduces a topological edge selection strategy to select targeted nodes for node injection, and then generate injected features through smooth feature optimization.
@@
  • FA-GNN [Hussain et al., 2022]: To the best of our knowledge, FA-GNN is the first work to conduct a fairness attack on GNNs. In detail, FA-GNN empirically discovers several edge injection strategies, which could impact the GNN fairness with slight utility compromise.
  • FATE [Kang et al., 2023]: FATE is a meta-learning-based fairness attack framework for GNNs. To be concrete, FATE formulates the fairness attacks as a bi-level optimization problem, where the lower-level optimization guarantees the deceptiveness of an attack while the upper-level optimization is designed to maximize the bias functions.
  • G-FairAttack [Zhang et al., 2024]: As a poisoning fairness attack, G-FairAttack consists of two modules, including a surrogate loss function and following constrained optimization for deceptiveness. Like FATE and FA-GNN, G-FairAttack also belongs to graph modification attacks, i.e., the original link structure between existing nodes will be modified during the attack.
@@ H.1 In-depth analysis of the poisoned graph
Structural analysis. Similar to prior work [Bojchevski et al., 2018, Sun et al., 2020], we investigate several key graph characteristics in this section, including Gini Coefficient, Assortativity, Power Law Exponent, Triangle Count, Relative Edge Distribution Entropy and Characteristic Path Length, whose definitions and implementations can be found here [^8] https://github.com/danielzuegner/netgan/tree/master. The graph statistics under different perturbation rates are shown in Table A2. It can be concluded that: (1) Thanks to the low perturbation rate required by NIFA, the poisoned graphs share quite similar characteristics with clean graphs. Under most cases, the relative rate of change $|\Delta|$ is smaller than 1%, especially when the perturbation rate is small. (2) With the increase of perturbation rate, the poisoning attack becomes more obvious, which can be verified by the increased $|\Delta|$ on all graph structure statistics. This observation further supports the necessity of requesting a low perturbation rate for an attack. (3) Several key statistics showcase a consistent trend across the three datasets as the perturbation rate increases. For example, with the increase in attack intensity, the degree assortativity consistently decreases on three datasets, indicating that there are more connections between nodes with significantly different degrees due to the injection of malicious nodes. Similar observations can also be found in key statistics like Triangle Count, Gini Coefficient and Characteristic Path Length, which implies that these statistics can be potentially utilized for fairness attack defense and auditing.
@@
Feature analysis. Besides graph structural analysis, we also conduct experiments to analyze the nodes’ features of the poisoned graph. In detail, after conducting fairness attacks through NIFA, for both labeled nodes in the real graph and injected nodes, we illustrate their features’ visualization results based on t-SNE [Van der Maaten and Hinton, 2008] in Figure A1. Note that, since the number of injected nodes is much smaller than that of the original labeled nodes, we slightly increase the scatter size of injected nodes for better visualization.
@@
H.2 Ablation study
Figure A2: Ablation study of each module in NIFA
The results are reported in Figure A2, where we can have the following observations. Firstly, after removing the uncertainty-maximization principle (NIFA-U), the fairness attack performance consistently decreases on three datasets. This is expected since the concept of uncertainty helps NIFA find more vulnerable nodes, thus improving the attack effectiveness. Secondly, after removing the homophily-increase principle (NIFA-H), the attack performance drops obviously, which verifies the homophily-ratio is crucial in GNN fairness. Finally, without iterative training during feature optimization (NIFA-I), the attack performance decreases slightly on all datasets. The main reason is that the iterative training strategy could help NIFA to have better robustness to dynamic victim models.
@@ H.3 Hyper-parameter analysis
H.4 Alternative analysis in node selection
During the node injection process, NIFA proposes an uncertainty-maximization principle and selects target nodes with the highest model uncertainty score. In fact, besides estimating model uncertainty, there are other alternative methods for finding vulnerable nodes. For example, TDGIA [Zou et al., 2021] introduces the concept of ”topological vulnerability”, and selects nodes with low degrees as target nodes. In this part, we also conduct experiments with selecting target nodes with the lowest degree as a variant of NIFA. The victim model is GCN.
Pokec-z Pokec-n DBLP
Clean 71.22 7.13 5.10 70.92 0.88 2.44 95.88 3.84 1.91
Degree 70.50 15.76 14.01 69.77 12.39 11.93 94.72 6.30 15.10
Uncertainty 70.50 17.36 15.59 70.12 10.10 9.85 93.37 13.49 20.33
H.5 Efficiency analysis
@@
To evaluate the efficiency and scalability of NIFA, we would like to compare the attack time cost and memory cost between NIFA and two competitive baselines including FATE [Kang et al., 2023] and G-FairAttack [Zhang et al., 2024]. To be fair, the attack budgets for the three models are set to be the same in advance. We report their memory cost and time cost for finishing poisoning attacks on Pokec-z, Pokec-n and DBLP in Table A4. The environment configurations for our experiments are introduced in Section G. It can be seen that, NIFA can be successfully deployed on three datasets with acceptable time cost and memory cost. In contrast, both FATE and G-FairAttack face scalability issues, especially when facing large graphs such as Pokec-z and Pokec-n. As shown in Table 1, both Pokec-z and Pokec-n have much more edges compared with DBLP, which causes FATE to report OOM errors and G-FairAttack to fail to finish the attack within 72 hours.

### v2

@@ Appendix D Implementation algorithm
Due to space limitation, here we provide the complete training process of NIFA in Algorithm 1. The training process and evaluation process are also literally described in Section 4.4.
@@ Appendix E Additional descriptions on victim models
  • GCN [10]: Borrowing the concept of convolution from the computer vision domain, GCN employs convolution operation on the graph from a spectral perspective to learn the node embeddings.
  • GraphSAGE [14]: Given the potential neighborhood explosion issues in GCN, GraphSAGE samples a fixed number of neighbors at each layer during neighborhood aggregation, which greatly improves the training efficiency.
  • APPNP [15]: Inspired by PageRank, APPNP decouples the prediction and propagation in the training process, which resolves inherent limited-range issues in message-passing models without introducing any additional parameters.
  • SGC [39]: SGC empirically finds the redundancy of non-linear activation function, and achieves comparable performance with much higher efficiency.
  • FairGNN [5]: Through proposing a sensitive attribute estimator and an adversarial learning module, FairGNN maintains high classification accuracy while reducing unfairness in scenarios with limited sensitive attribute information.
  • FairVGNN [38]: FairVGNN discovers the leakage of sensitive information during information propagation of GNN models, and generates fair node features by automatically identifying and masking sensitive-correlated features.
  • FairSIN [40]: Instead of filtering out sensitive-related information for fairness, FairSIN deploys a novel sensitive information neutralization mechanism. Specifically, FairSIN will learn to introduce additional fairness facilitating features (F3) during message propagation to neutralize sensitive information while providing more non-sensitive information.
@@ Appendix F Additional descriptions on baselines
  • AFGSM [36]: As a node injection-based attack, AFGSM designs an approximation strategy to linearize the victim model and then generates adversarial perturbation efficiently. In general, AFGSM is scalable to much larger graphs.
  • TDGIA [47]: Aiming at attacking the model performance, TDGIA first introduces a topological edge selection strategy to select targeted nodes for node injection, and then generate injected features through smooth feature optimization.
@@
  • FA-GNN [11]: To the best of our knowledge, FA-GNN is the first work to conduct a fairness attack on GNNs. In detail, FA-GNN empirically discovers several edge injection strategies, which could impact the GNN fairness with slight utility compromise.
  • FATE [13]: FATE is a meta-learning-based fairness attack framework for GNNs. To be concrete, FATE formulates the fairness attacks as a bi-level optimization problem, where the lower-level optimization guarantees the deceptiveness of an attack while the upper-level optimization is designed to maximize the bias functions.
  • G-FairAttack [41]: As a poisoning fairness attack, G-FairAttack consists of two modules, including a surrogate loss function and following constrained optimization for deceptiveness. Like FATE and FA-GNN, G-FairAttack also belongs to graph modification attacks, i.e., the original link structure between existing nodes will be modified during the attack.
@@ H.1 In-depth analysis of the poisoned graph
Structural analysis. Similar to prior work [2, 32], we investigate several key graph characteristics in this section, including Gini Coefficient, Assortativity, Power Law Exponent, Triangle Count, Relative Edge Distribution Entropy and Characteristic Path Length, whose definitions and implementations can be found here [^8] https://github.com/danielzuegner/netgan/tree/master. The graph statistics under different perturbation rates are shown in Table A2. It can be concluded that: (1) Thanks to the low perturbation rate required by NIFA, the poisoned graphs share quite similar characteristics with clean graphs. Under most cases, the relative rate of change $|\Delta|$ is smaller than 1%, especially when the perturbation rate is small. (2) With the increase of perturbation rate, the poisoning attack becomes more obvious, which can be verified by the increased $|\Delta|$ on all graph structure statistics. This observation further supports the necessity of requesting a low perturbation rate for an attack. (3) Several key statistics showcase a consistent trend across the three datasets as the perturbation rate increases. For example, with the increase in attack intensity, the degree assortativity consistently decreases on three datasets, indicating that there are more connections between nodes with significantly different degrees due to the injection of malicious nodes. Similar observations can also be found in key statistics like Triangle Count, Gini Coefficient and Characteristic Path Length, which implies that these statistics can be potentially utilized for fairness attack defense and auditing.
@@
Feature analysis. Besides graph structural analysis, we also conduct experiments to analyze the nodes’ features of the poisoned graph. In detail, after conducting fairness attacks through NIFA, for both labeled nodes in the real graph and injected nodes, we illustrate their features’ visualization results based on t-SNE [35] in Figure A1. Note that, since the number of injected nodes is much smaller than that of the original labeled nodes, we slightly increase the scatter size of injected nodes for better visualization.
@@
H.2 Alternative analysis in node selection
Pokec-z Pokec-n DBLP
Clean 71.22 7.13 5.10 70.92 0.88 2.44 95.88 3.84 1.91
Degree 70.50 15.76 14.01 69.77 12.39 11.93 94.72 6.30 15.10
Uncertainty 70.50 17.36 15.59 70.12 10.10 9.85 93.37 13.49 20.33
During the node injection process, NIFA proposes an uncertainty-maximization principle and selects target nodes with the highest model uncertainty score. In fact, besides estimating model uncertainty, there are other alternative methods for finding vulnerable nodes. For example, TDGIA [47] introduces the concept of ”topological vulnerability”, and selects nodes with low degrees as target nodes. In this part, we also conduct experiments with selecting target nodes with the lowest degree as a variant of NIFA. The victim model is GCN.
@@ H.3 Hyper-parameter analysis
H.4 Efficiency analysis
@@
To evaluate the efficiency and scalability of NIFA, we would like to compare the attack time cost and memory cost between NIFA and two competitive baselines including FATE [13] and G-FairAttack [41]. To be fair, the attack budgets for the three models are set to be the same in advance. We report their memory cost and time cost for finishing poisoning attacks on Pokec-z, Pokec-n and DBLP in Table A4. The environment configurations for our experiments are introduced in Section G. It can be seen that, NIFA can be successfully deployed on three datasets with acceptable time cost and memory cost. In contrast, both FATE and G-FairAttack face scalability issues, especially when facing large graphs such as Pokec-z and Pokec-n. As shown in Table 1, both Pokec-z and Pokec-n have much more edges compared with DBLP, which causes FATE to report OOM errors and G-FairAttack to fail to finish the attack within 72 hours.
H.5 Robustness to defense strategies
Besides the defense discussions in Section 6, we analyze the effectiveness of conventional defense strategies in this section. In detail, two classic GNN defense models – GNNGuard [44] and ElasticGNN [18] are deployed to test their defense capabilities against NIFA. It is worth noting that, both defense models are originally designed for the attacks on prediction utility. We conduct experiments on three datasets, and the victim model is GCN as default. As shown in Table A5, both defense models only maintain the utility performance and failed to fully eliminate the impact of fairness attacks. For example, compared with the performance after the attack, although ElasticGNN reduces the $\Delta_{SP}$ from 17.36% to 13.53%, the fairness is still worse than that before the attack – 7.13%. We believe that the main reason behind such observation is that all these defense methods are designed for utility-targetted attacks, whose objectives are totally different from fairness attacks like NIFA. Such observations also indicate that more effective defense mechanisms are still in demand for fairness attacks on GNNs.
Table A5: The performance of defense strategies against NIFA.
Pokec-z Pokec-n DBLP
Acc $\Delta_{SP}$ $\Delta_{EO}$ Acc $\Delta_{SP}$ $\Delta_{EO}$ Acc $\Delta_{SP}$ $\Delta_{EO}$
Clean 71.22±0.28 7.13±1.21 5.10±1.28 70.92±0.66 0.88±0.62 2.44±1.37 95.88±1.61 3.84±0.34 1.91±0.75
NIFA 70.50±0.30 17.36±1.16 15.59±1.08 70.12±0.37 10.10±2.80 9.85±1.97 93.37±0.28 13.49±2.83 20.33±3.82
GNNGuard 71.50±0.35 18.13±2.31 16.32±1.44 70.55±0.43 13.82±1.76 13.38±1.42 95.73±0.28 7.42±1.82 14.94±1.58
ElasticGNN 71.36±0.20 13.53±0.92 11.36±1.23 70.32±0.28 6.90±1.02 6.44±1.32 94.22±0.37 9.76±1.65 17.04±2.01
H.6 Performance with limited training ratio
Table A6: The performance of NIFA with more limited training nodes. The victim model is GCN.
Pokec-z Pokec-n
Acc $\Delta_{SP}$ $\Delta_{EO}$ Acc $\Delta_{SP}$ $\Delta_{EO}$
25% before 71.47±0.59 6.26±1.63 4.23±1.55 70.16±0.45 1.03±0.70 2.87±0.53
after 70.48±0.41 15.16±1.64 13.18±1.77 69.49±0.31 9.79±2.02 9.52±1.84
10% before 70.40±0.28 6.29±2.77 4.63±2.43 70.25±0.57 2.56±1.32 3.72±0.32
after 70.14±0.26 15.26±3.02 13.36±3.06 69.25±0.62 11.94±1.78 12.00±1.75
5% before 70.03±0.54 4.90±1.98 3.75±1.01 68.61±1.15 3.79±2.30 4.21±1.94
after 69.49±0.65 14.35±2.03 12.42±2.04 68.05±1.63 10.31±5.68 10.60±5.04
As introduced in Section 5.1, our datasets are all collected from the previous work – FA-GNN [11], and our train / val / test ratios are set to be consistent with its settings, which is 50% / 25% / 25%, respectively. However, in more realistic scenarios the percentage of labeled training nodes might be significantly smaller due to the heavy cost of annotation. To evaluate the effectiveness of NIFA under such settings, we decrease the training ratio from 50% to 25%, 10% and 5%, respectively. The attack performance of NIFA on Pokec-z and Pokec-n is shown in Table A6. It can be seen that, even with much fewer labeled training nodes, NIFA still consistently demonstrates promising attack performance.